OpenStudy (anonymous):

a feild is to be fertilized at a cost of$0.08 per square yard. The rectangular part of the field is 95 yards long and the diameter of each semicircle is 49 yds. Find the cost of fertilizing the field.

OpenStudy (anonymous):

Find the area of the field if the rectangular part is 95 yds long and the diameter of the semicircles is 49 yds, the area is: 95*49 that's just for rectangular area now add those two semicircles and find the whole area of a circle A = pi*r^2 = pi * (49/2)^2 Total area: 95*49+pi*(49/2)^2 Total cost: $0.08*(95*49+pi*(49/2)^2)

OpenStudy (anonymous):

A. $3,92387 B. $420.42 C.$975.53 D.$523.18

OpenStudy (anonymous):

Is it: